# Efficient Deep Spiking Multi-Layer Perceptrons with Multiplication-Free Inference

### v1

@@ IV-D Ablation Study
In this section we further study the influence of different configurations of skip connections on the performance of the network. As shown in Figure 4, we distinguish different skip connections, i.e patch block to token block (PT), patch block to channel block (PC), token block to channel block (TC) and channel block to the next token block (CT), with corresponding indices and colors. The skip connection from the patch block to the first token block is set as default. To shorten the simulation time, we set the initial channel number to 60 and directly train networks of different skip connection configurations from scratch on the CIFAR10 dataset for 100 epochs. The results are collected in Table IV. Networks without skip connections from the initial patch block to sequential token and channel blocks perform much worse than the others. This could be due to long range skip connections alleviate the gradient vanish problem of deep SNNs during training. Our spiking MLP-mixer adopts the optimal skip connection configuration of (PT, PC, TC).
@@
Table IV: Network performance under different skip connection configurations on CIFAR10.
@@
IV-E Network Sparsity and Computational Cost
Figure 5: Mean network sparsity of spiking MLP-SPE-T on the ImagNet-1K test set. We distinguish each stage with different colors.
SNNs are recognized for their superior energy efficiency compared to dense-computing ANNs, a quality that primarily stems from the inherent event-driven and sparse computing characteristics of SNNs. Additionally, the MFI principle facilitates an entirely addition-based operation for SNNs, thus further diminishing their potential energy usage in contrast to ANNs that necessitate real-valued matrix multiplications. In this section, we assess the sparsity and computation cost of the spiking MLP-SPE-T model, contrasting it with the spiking ResNet-34.
Figure 5 portrays the average sparsity of each spiking MLP-Mixer on the ImageNet-1K test set. Our model demonstrates diverse sparse activity across different layers, yielding an average network sparsity of 0.14. Following the methodologies established by [15, 30], we compute the number of addition operations in the SNN using the formula $sTA$, where $s$ represents the average sparsity (the total count of spikes divided by the total feature map size), $T$ indicates the simulation time step, and $A$ signifies the number of additions.
The findings are presented in Table V. The least amount of multiplication arises from the input layer, which receives floating-value images. To estimate the network’s energy consumption, we refer to the study by [45] on the $45\mathrm{nm}$ CMOS technology, a method also employed by previous works [15, 30]. According to this estimate, a single addition operation in an SNN costs $0.9\mathrm{pJ}$, whereas a multiply-accumulate (MAC) operation in ANNs consumes $4.6\mathrm{pJ}$. Notwithstanding their comparable model capacities, our network’s computation cost is nearly half of that of the spiking ResNet-34.
Table V: The computation and estimated energy cost.
@@ IV-F Visualization
Figure 6: Axial sampling weights from token blocks across different stages after training on ImageNet-1K. Figures (a-c) show randomly selected token weights plotted in a $H\times H$ or $W\times W$ 2D matrix. 1, 2, and 4 weight matrices from stage 1, 2, and 3 are displayed, respectively. The diagonal distribution of values in the early stages indicates local sampling. Figure (d) represents weights of several neurons from the same token block in stage 2. For the sake of simplicity, we only plot a few neurons here; however, it should be noted that this ’M’ shaped weight kernel is a universal finding in the network’s early stages.
@@ V Conclusion
Our work underscores the importance of integrating global and local learning for optimal SNN architecture design. Interestingly, the trained receptive fields of our network bear a striking resemblance to those of cells in the cortex. Future comparative analysis with ANNs could potentially yield insightful results.

### v3

@@ IV-D Ablation Study
IV-D1 Normalization Methods
Table IV: Comparisons of different normalization methods on ImageNet-1K (T denotes simulation length)
[…]
MLP-SPE-T Batch Norm 25M 4 $66.39$
MLP-SPE-T Layer Norm 25M 4 $65.94$
MLP-SPE-T TEBN 25M 4 $62.33$
MLP-SPE-T tdBN 25M 4 $60.25$
We conduct additional experiments focused on replacing the BN operation within the MLP-Block by other normalization methods including the original LN [56] in the MLP-Mixer and two recent temporal normalization methods specifically proposed for SNNs, i.e., tdBN [45] and TEBN [76]. These experiments are conducted on the ImageNet-1k dataset, maintaining consistency with the hyperparameters used in the ImageNet experiments. The results of these experiments are presented in Table IV. Our findings indicate that BN suggests more beneficial for accuracy improvement compared to other normalization methods. Additionally, employing BN allows the parameters of the latter to integrate with the linear projection weights during inference, aligning with the MFI principle.
IV-D2 Skip Connection
We further study the influence of different configurations of skip connections on the performance of the network. As shown in Figure 4, we distinguish different skip connections, i.e patch block to token block (PT), patch block to channel block (PC), token block to channel block (TC) and channel block to the next token block (CT), with corresponding indices and colors. The skip connection from the patch block to the first token block is set as default. To shorten the simulation time, we set the initial channel number to 60 and directly train networks of different skip connection configurations from scratch on the CIFAR10 dataset for 100 epochs. The results are collected in Table LABEL:tab:skip. Networks without skip connections from the initial patch block to sequential token and channel blocks perform much worse than the others. This could be due to long range skip connections alleviating the gradient vanish problem of deep SNNs during training. Our spiking MLP-mixer adopts the optimal skip connection configuration of (PT, PC, TC).
@@ IV-D2 Skip Connection
Table V: Network performance under different skip connection configurations on CIFAR10
@@
IV-E Network Spiking Rate and Computational Cost
Figure 5: Mean network spiking rate of MLP-SPE-T on the ImagNet-1K test set. We distinguish each stage with different colors.
SNNs are recognized for their superior energy efficiency compared to dense-computing ANNs, a quality that primarily stems from the inherent event-driven and sparse computing characteristics of SNNs. Additionally, the MFI principle facilitates an entirely addition-based operation for SNNs, thus further diminishing their potential energy usage in contrast to ANNs that necessitate real-valued matrix multiplications. In this section, we assess the spiking rate and computation cost of the spiking MLP-SPE-T model, contrasting it with the spiking ResNet-34.
Figure 5 portrays the average spiking rate of each spiking MLP-Mixer on the ImageNet-1K test set. Our model demonstrates diverse sparse activity across different layers, yielding an average network spiking rate of 0.14. Following the methodologies established by [15, 30], we compute the number of addition operations in the SNN using the formula $sTA$, where $s$ represents the average spiking rate (the total count of spikes divided by the total feature map size), $T$ indicates the simulation time step, and $A$ signifies the number of additions.
The findings are presented in Table VI. The least amount of multiplication arises from the input layer, which receives floating-value images. To estimate the network’s energy consumption, we refer to the study by [49] on the $45\mathrm{nm}$ CMOS technology, a method also employed by previous works [15, 30]. According to this estimate, a single addition operation in an SNN costs $0.9\mathrm{pJ}$, whereas a multiply-accumulate (MAC) operation in ANNs consumes $4.6\mathrm{pJ}$. Notwithstanding their comparable model capacities, our network’s computation cost is nearly half of that of the spiking ResNet-34. The SparseMLP [36] in the ANN domain and the SpikFormer [72], which employs a spike-based attention mechanism, achieve significantly higher accuracy but consume over ten times the energy compared to our model.
Table VI: The computation and estimated energy cost
@@ IV-F Visualization
Figure 6: Axial sampling weights from token blocks across different stages after training on ImageNet-1K. Figures (a-c) show randomly selected token weights plotted in a $H\times H$ or $W\times W$ 2D matrix. 1, 2, and 4 weight matrices from stage 1, 2, and 3 are displayed, respectively. The diagonal distribution of values in the early stages indicates local sampling. Figure (d) represents weights of several neurons from the same token block in stage 2. For the sake of simplicity, we only plot a few neurons here, however, it should be noted that this ’M’ shaped weight kernel is a universal finding in the network’s early stages.
@@ V Conclusion
Our work underscores the importance of integrating global and local learning for optimal SNN architecture design. Interestingly, the trained receptive fields of our network bear a striking resemblance to those of cells in the cortex. Future comparative analysis with ANNs could potentially yield insightful results. In terms of accuracy, however, our SNN-based MLPs are still suboptimal compared with MLP architectures in the ANN domain. Considering this is an initial effort in adapting SNNs to MLP frameworks, future research into more effective structures and spike-based mechanisms, such as adaptive thresholds, is likely to enhance network performance. Furthermore, recent works of SNNs integrated with attention mechanisms have been applied to challenging language tasks [78, 79]. With the ever growing demand on computational efficient architectures for these tasks, the proposed spike-based MLP structure could be integrated as a light weight module in the encoder network.